# Forge: Foundational Optimization Representations from Graph Embeddings

### v2

@@ 3 Forge: Unsupervised Approach to Learn MIP Embeddings
Local representations, where each node in the bipartite graph is assigned a discrete code. Each code is further mapped to a codeword which becomes the variable and/or constraint embedding.
Global representations, where we leverage the distribution of codes for the given MIP instance. As shown in Figure 2, each MIP instance is represented with an embedding of size $vocabulary=\lvert codebook\rvert$ where each value denotes the frequency of codes present in the instance. Figure 2 shows a codebook of size 5 (for brevity) and the resulting MIP embedding $\vec{emb}=[3,2,3,2,0]$.
Note that this form of unsupervised pre-training aims at learning the structure of MIP instances and is different from methods which aim to cast the discrete objective function of a MIP instance as a differentiable one and learn the solution using gradient descent in an end to end manner (Sanokowski et al., 2024; Karalias & Loukas, 2020; Bu et al., 2024).
Next, we start with an initial investigation of the effectiveness of Forge embeddings when clustering unseen instances across various problem domains (§4), and then explore how they can enhance solving MIP instances via warm-starts and cut-generation(§5).
4 Initial Analysis of Forge Embeddings
Given the absence of methods in the literature that can provide general-purpose MIP embeddings at the instance level, with the exception of earlier works that depend on creating hand-crafted features to classify MIP instances (e.g., (Kadioglu et al., 2010)), we start our initial analysis with a comparison of Forge embeddings against two (ablation) baselines when clustering unseen instances.
Our goal is to investigate both the accuracy of the clustering (quantitative) as well as visual inspection (qualitative). For this purpose, we train Forge on a set of MIP instances from MIPLIB (Gleixner et al., 2021), and test it on unseen instances from Distributional MIPLIB (D-MIPLIB) (Huang et al., 2024). While MIPLIB is a mixed dataset, D-MIPLIB is categorized into different problem classes. This serves as ground truth label when we treat each problem class as a cluster to evaluate our embeddings.
Figure 3: 2-D PaCMAP visualization of D-MIPLIB instances using MIP embeddings from (a) the mean readout of node embeddings from the GNN in Forge, (b) the two-hop label propagation of input node features, and (c) the distribution of discrete codes produced by Forge. Normalized Mutual Information (NMI), averaged over 10 runs of $k$-means clustering, is reported for each method.
Training: We use 600 instances from MIPLIB, sorted by size to ensure the resulting bipartite graphs fit on GPU memory. We further generate two more instances from each MIPLIB instance by randomly dropping 5% and 10% of constraints. Note that dropping constraints does not impact feasibility of the instance. In total, we obtain 1,800 MIP instances to train Forge in an unsupervised fashion as described earlier (§3). We use two GraphSage layers with 1024 dimensions for the hidden GNN layers ($d$) and a codebookcodebook with 5000 codes (i.e., the size of the vocabulary).
Testing: We evaluate Forge on 1,050 instances from D-MIPLIB organized into 21 problem type-difficulty pairs. These are distributed as Set Cover (easy, medium, hard), Maximum Independent Set (easy, medium), Minimum Vertex Cover (easy, medium, hard), Generalized Independent Set (easy, medium, hard, very-hard, very-hard2, ext-hard), Combinatorial Auction (very-easy, easy, medium, very-hard, very-hard2), Item Placement (very-hard) and Maritime Inventory Routing Problem (medium), thereby covering a broad spectrum of problem types, sizes and difficulty levels.
MIP Embeddings: These 1,050 instances are passed through Forge to generate one embedding vector per MIP instance using the distribution of discrete codes as shown in Figure 2. For comparison, we consider two other alternatives. As a baseline approach, we use the Mean Readout utilizing the GNN embeddings of the trained Forge model. The Mean Readout method generates the MIP embedding of an instance by averaging all node features from the GNN hidden layer representation from Figure 1(B). This baseline also behaves as an ablation for Forge without vector quantization. Given the weakness of GNN embeddings in capturing global structure (Feng et al., 2025), we expect this method to perform poorly. Alternatively, starting from the 10-dimensional static node features of the bipartite graph (Fig. 1(A)), we perform two-hop label propagation (Zhu & Ghahramani, 2002) and average the resulting node vectors. Note that global, static instance descriptors (e.g., (Kadioglu et al., 2010)) cannot be used since within D-MIPLIB, every instance in a category shares identical coarse statistics (number of variables, constraints, etc.), rendering them non-informative.
Clustering Visualization: Figure 3 visualizes these embeddings vectors, projected into two dimensions using PaCMAP (Wang et al., 2021). Each dot represents an instance colored by its problem type. As expected, the Mean Readout method looses the global structure leading to an arbitrary clustering. The Label Propagation performs much better capturing some structure, while some problem domains and sizes are mixed in the clustering. Forge embeddings performs the best where each problem type, such as Set Cover, Combinatorial Auction etc., are separated cleanly. Moreover, within each problem domain, the difference in problem difficulty is also observed from the easy to medium and hard categories. Recall again that our Forge model is not trained on these instances. We are able to train on MIPLIB instances and capture the structure of unseen D-MIPLIB instances.
Clustering Accuracy: For quantitative evaluation, we run k-means with 21 clusters, expecting one cluster for each problem type. We calculate the normalized mutual information score (NMI) between then ground truth problem type assignment and the k-means clusters generated by each method. The NMI scores are averaged over 10 runs of k-means clustering. If the predicted clusters are identical to the original problem types, then NMI score would be 1.0, and 0.0 otherwise.
As qualitatively observed from the 2D visualization, the Mean Readout performs poorly with an NMI score of 0.087. This is only slightly better than random cluster predictions with an NMI score of $1/21=0.047$. This is likely due to over-smoothing due to averaging dense GNN embedding vectors across a large number of constraints and variables. In contrast, the Label Propagation performs better with an NMI score of 0.790. This method operates directly on the sparse input features of the bipartite graph, thereby avoiding the over-smoothing problem encountered in deep GNN layers. The best performance is achieved by Forge with an NMI score of 0.843. Interestingly, the mean readout embeddings, which perform very poorly, are generated from the same trained GNN within Forge. By utilizing the distribution of the discrete codes of constraints and variables, Forge circumvents the over-smoothing issue and captures the global structure of unseen MIP instances.
5 Experiments
Having demonstrated that Forge embeddings reliably cluster unseen instances from diverse problem classes, we now shift to supervised evaluations more directly aimed at improving MIP solving. Accordingly, we design the next suite of experiments on:
  • Downstream tasks that are commonly applied in the literature to enable us with a fair comparison and provide utility to enhance MIP solving.
  • Downstream tasks that are radically different from each other, where we use the same Forge model, to validate its general applicability across tasks, problems, and sizes.
  • Downstream tasks that are agnostic to the underlying MIP solver, i.e., we do not depend on internal access to specific solver procedures (e.g., within the branch-and-bound tree).
We therefore study two different downstream tasks: predicting the integrality gap of a given MIP instance, as also used in (Li et al., 2025), and predicting variables for warm-starts, as also used in (Han et al., 2023), in Sections § 5.2 and § 5.3 respectively. Both tasks serve as primal heuristics to enhance MIP solving by obtaining better solutions faster. The integrality gap is used to generate a pseudo-cut added to the original problem formulation to tighten its bound. The warm-start variables are used to provide hints to the solver to start the solving process from a promising initial assignment.
A critical aspect of these experiments is that we use the same pre-trained Forge model to obtain general-purpose MIP embeddings that are then fine-tuned on a small number of labeled data to learn prediction heads for completely different tasks (as shown in Figure 4(A)). An analogy from NLP would be to generate pre-trained word embeddings using a foundational model, and then fine-tune prediction heads for entity extraction in a specific domain, e.g., finance, using a small set of labeled data. Our goal is to replicate the success of foundational models in NLP and CV in the context CO.
We begin with an overview of our training setup to create our pre-trained Forge model and then discuss its applications to predicting the integrality gap for pseudo-cut generation and predicting warm-start variables for guiding the search.
5.1 Training Setup for Foundational Forge Model
The setup for training Forge is identical to the setting in the initial clustering analysis, with two GraphSage layers with $d=1,024$ dimensions and $k=5000$ codes. The only difference is that we pre-train on the structure of the 1,800 MIPLIB instances as well as the 1,050 D-MIPLIB instances. In total, Forge is trained on a corpus of 2,850 MIP instances.
@@ 5.1 Training Setup for Foundational Forge Model
For these experiments, we train on the ml.g5.xlarge AWS instance with a GPU with 24 GB of memory. Inference experiments were run on the ml.c5.12xlarge instance with 48 cores and 96 GB of RAM. To ensure consistency and fairness, all experiments were executed with Gurobi, a state-of-the-art commercial MIP solver (Gurobi Optimization, LLC, 2024), restricted to a single thread and a time limit of 3600 seconds. Unsupervised pre-training and integrality gap fine-tuning are run for 10 epochs with a learning rate of $10^{-4}$, while the warm-start prediction task is trained for 25 epochs using a learning rate of $10^{-5}$.
5.2 Integrality Gap Prediction & Pseudo-Cut Generation
Integrality Gap Prediction: As covered earlier (§ 2), the integrality gap measures the ratio between the optimal solution of the LP relaxation and the optimal solution of the original integer program. Intuitively, this gap quantifies the quality of the approximation offered by the LP relaxation. A smaller gap means the LP relaxation is a good approximation and is close to the value of the best integral solution. Here, we are interested in predicting the integrality gap of a given MIP instance, without solving for its optimal integral solution value, as also studied in (Li et al., 2025).
Pseudo-Cut Generation: If we can predict the integrality gap, then we can generate a pseudo-cut and add it as an additional constraint to the model to immediately bound the optimal objective value from the initial LP relaxation. Note that, this cut is not guaranteed to be a valid cut at all times, hence it is a pseudo-cut. If the integrality gap prediction is incorrect, it risks over (or under) estimating the best objective value. This makes integrality gap prediction a challenging problem.
Figure 4: (A) The supervised fine tuning flow for Forge. Integrality gap predictions are generated by using a mean readout and warm start probabilities are generated for each variable using a sigmoid function. (B) Distribution of integrality gaps for different problem type-difficulty pairs. (C) Procedure for constructing triplets for the triplet loss for warm start prediction. Each MIP instance is solved to obtain the top 5 solutions. Nodes appearing in the same number of solutions are paired as anchor and positive, while nodes absent from all solutions are designated as negatives.
Training Instances: We do not expect this task to generalize between problem classes and/or sets of varying complexities of a given problem class. For instance, there is no reason for an LP gap of 70% to be the same between easy vs. hard SC instances. Figure 4(B) shows the distribution of integrality gap across different problem types and difficulties. Note that the integrality gap can be anywhere from 5% to 95% with wide distributions in between. As such, there is no magic constant that one could use at all times heuristically, which makes integrality gap prediction a deliberate learning task. For training, we consider CA (very-easy, easy, medium), SC (easy, medium, hard), and GIS (easy, medium, hard) with 50 instances for each. In total, we obtain a total of 450 training instances. Note that this is a considerably smaller training set than initially used for training Forge embeddings.
Label Collection: To generate training labels, each training instance is solved using Gurobi with a time limit of 120 seconds. Note that this does not require solving the instance to optimality which can be computationally challenging. The numeric label is defined as the ratio (or gap) between the integer program solution at the time-out value and the LP relaxation solution. As mentioned, overestimating the integrality gap (or underestimating for maximization problems) can lead to suboptimal solutions, as the solver may terminate prematurely. In contrast, underestimating the gap is generally acceptable, as it still facilitates faster solve times without compromising solution quality. This asymmetry is accounted for by adopting a conservative labeling strategy by setting a 120-second timeout when collecting labels. This often results in an underestimation of the true integrality gap, particularly for more challenging instances. By doing so, we reduce the risk of the model overestimating the gap, thereby improving the reliability of the predicted cuts.
Supervised Fine-Tuning: Given the labeled data, a dense prediction head is added to the pre-trained Forge model that takes as input the codewords assigned to each node and outputs a real number (Figure 4(A)). The mean readout across all nodes is used as the predicted integrality gap. As a regression task, this is trained with the mean absolute error loss in an end-to-end manner.
5.2.1 Numerical Results: Integrality Gap Prediction
Test Instances & Setup: We start with the pre-trained Forge model, fine-tuned with a prediction head using the labeled data, and predict the integrality gap of 50 ‘very-hard’ instances each of CA, SC, GIS, and MVC. Given the prediction, a pseudo-cut is generated by adjusting the initial LP relaxation objective - by increasing it for minimization problems or decreasing it for maximization problems. This pseudo-cut is incorporated into the original formulation as an additional constraint, enforcing that the integral objective must exceed (or fall below) the computed pseudo-cut value.
Prediction Accuracy: We measure the deviation in terms of mean absolute error between the true integrality gap and the gap predicted by Forge. On the ‘very-hard’ instances, Forge achieves a deviation of $15.42\%$, $13.55\%$, $12.03\%$ and $19.077\%$ for CA, SC, GIS and MVC instances respectively. Note that MVC instances were not part of the train set.
While predicting integrality gaps may appear straightforward, training Forge from scratch for this task led to a $\approx 33$% increase on average in mean absolute error across all problem types compared to fine-tuning a pre-trained model. This highlights the importance of unsupervised pre-training in capturing transferable structural patterns across diverse MIP instances.
Comparison with the Commercial MIP Solver: We compare the default Gurobi solver on these ‘very-hard‘ instances with and without our predicted pseudo-cut. Figure 5 shows the primal gap averaged over these 50 ‘very-hard’ instances of each type with a time limit of 3600 seconds. Without exception, across all problem types, the use of pseudo-cuts generated by Forge embeddings consistently results in lower primal gaps over time. The gains range from 30% to 85%. The default solver attains immediate improvements early on in the search and our pseudo-cuts make these gains more pronounced immediately. Recall that this only needed 50 instances from each problem type.
Figure 5: Gurobi vs. Gurobi + Pseudo-Cuts generated by Forge. Each subplot shows the primal gap reduction averaged across 50 ‘very-hard’ instances per problem type (lower is better).
Comparison with SOTA ML Methods: To further evaluate generalization across problem types and sizes, we compare against the setup described in  Li et al. (2025), where a GNN is trained on 38,256 instances from 643 generated problem types and tested on 11,584 instances spanning 157 problem types. We use Forge as is, not using any additional training instances apart from 450 D-MIPLIB instances, and test on 17,500 previously unseen instances spanning 400 generated problem types. Forge achieves a mean deviation of 18.63% in integrality gap prediction, outperforming the reported 20.14% deviation by almost 10%.
5.3 Warm-Start Prediction & Guiding The Search
@@
Warm Start Prediction: Recall that Forge generates embedding vectors per instance and per variable/constraint. We fine tune the variable embeddings to predict which variables are likely to be in the optimal solution. These predictions are then passed into Gurobi as variable hints. A straightforward method of predicting which variables are likely to be part of the solution is to treat it as a binary classification problem and use binary cross entropy (BCE) loss. However, this poses challenges due to the large class imbalance where most variables are not part of the solution.
@@
To address this issue, we generate 5 feasible solutions for each instance. Variables in the instance are grouped by how many of the 5 solutions each variable was repeated in. A variable is considered a “negative” variable if and only if it appears in none of the 5 solutions. Next, we use triplet loss (Schroff et al., 2015) to fine tune the embeddings of the variables. The triplet loss is given by
@@
Training Instances: We collect 100 instances each from CA (easy, medium), SC (easy, medium, hard) and GIS (easy, medium) for a total of 700 training instances.
Label Collection: Each instance is solved using Gurobi and a solution pool of 5. Here, we set a larger time limit of 300 seconds since we would like to get closer to the optimal solution.
Supervised Fine Tuning: The triplet loss is used to fine tune the GNN within Forge to encourage meaningful separation in the embedding space. In parallel, a dense prediction head is added to Forge that takes as input the codewords assigned to each node (Figure 4(C)). The model is trained using both the triplet loss and binary cross entropy loss on the prediction head.
5.3.1 Numerical Results: Warm-Start Prediction
Test Instances & Setup: We start with the pre-trained Forge model, fine-tuned with a prediction head using the labeled data, and then predict warm start variables of 50 ‘medium’ instances each of Combinatorial Auction (CA), Set Cover (SC), Generalized Independent Set (GIS), and Minimum Vertex Cover (MVC). Note that MVC instances are not part of the train set.
To predict the warm start variables, we begin by generating a feasible solution using Gurobi with a 1-second timeout. The variables included in this solution are designated as seed variables. Since fine-tuning with triplet loss encourages nodes corresponding to solution variables to cluster in the embedding space, we identify additional variables whose embeddings lie within a fixed radius (set to 0.1) of the seed variables. Simultaneously, we select variables ranked in the top 5th percentile and bottom 10th percentile by the prediction head as positive and negative candidates, respectively. The intersection of variables identified by the embedding-based proximity search and the prediction head are then passed to Gurobi as solution hints.
Comparision with the Commercially MIP Solver: We compare the default Gurobi solver on these ‘medium‘ instances with and without our predicted warm start variables. Figure 6 shows primal gaps averaged over 50 instances for each of CA, SC, GIS and MVC using Gurobi versus Gurobi warm started with predicted variables with a time limit of 3600 seconds. The warm-started variants consistently achieve lower primal gaps and converge to optimal solutions significantly faster, demonstrating the practical utility of Forge in accelerating MIP solving.
@@ 5.3.1 Numerical Results: Warm-Start Prediction
Figure 6: Gurobi vs. Gurobi + Warm Starts generated by Forge. Each subplot shows the primal gap reduction averaged across 50 ‘medium’ instances per problem type (lower is better). Gurobi + Warm Starts also converges to the optimal solution significantly faster.
Figure 7: Performance comparison of three variants of PS-Gurobi: the original model, a version augmented with 64-dimensional random vectors, and a version augmented with Forge embeddings (reduced to 64 dimensions via PCA). Evaluation is conducted on 50 ‘medium’ instances each of CA and GIS problems. The Forge-augmented PS-Gurobi consistently outperforms the baseline, achieving average reductions in primal gap of 41.07% for CA and 50.51% for GISP.
Comparison with SOTA ML Methods: As an additional validation of the effectiveness of Forge embeddings, we augment PS-Gurobi (Han et al., 2023), a state-of-the-art ML-based method that already demonstrates strong performance relative to other learning-based approaches. Specifically, we augment the input node features for variables and constraints in PS-Gurobi by concatenating the unsupervised embeddings generated by the pre-trained Forge model.
Since Forge embeddings for each node are 1024-dimensional, we apply Principal Component Analysis (PCA) to reduce the dimensionality to 64 dimensions, thereby minimizing architectural modifications to the original model. To further ensure that any observed performance gains are attributed to the semantic content of the Forge embeddings rather than increased model capacity, we also conduct a control experiment by augmenting PS-Gurobi with 64-dimensional random vectors. This comparison allows us to isolate the contribution of the learned embeddings to overall performance. We evaluate this setup on the same 50 instances each of the CA and GIS problems - two problem types that are common to both Forge and PS-Gurobi. Results are shown in Figure 7 where PS-Gurobi + Forge consistently outperforms the standalone PS-Gurobi with the gain in primal gaps being 41.07% and 50.51% for CA and GIS respectively. Interestingly, PS-Gurobi with random vectors occasionally surpasses the original model, likely due to the significant increase in input dimensionality - from 4 and 6 features for constraints and variables, respectively, to 68 and 70 - resulting in a larger model capacity that can yield marginal performance gains.
6 Related Work

### v4

@@ A.2 Example of Instance Level MIP Embeddings
A.3 Additional Clustering Results on strIPlib
To further validate our clustering results from §4, where we pre-trained Forge on 1,800 MIPLIB instances and clustered instances from D-MIPLIB, we repeat the experiment using our MIPLIB-pre-trained Forge to cluster instances from strIPlib (Bastubbe et al., 2025). We select 50 instances from each of 10 previously unseen problem types: Bin Packing, Capacitated Warehouse Location, Cutpacking, General Assignment, Job Shop Scheduling, Map Labeling, Scheduling, Stochastic Server Location, Train Timetabling, and Vertex Coloring.
Figure 8 shows 2D clustering visualizations using PaCMAP. As in our previous study, these strIPlib instances were unseen by Forge, yet the model still clusters different problems cleanly. For example, all packing instances group together in the top-left, while warehouse and server location instances cluster in the bottom-left. Other interesting patterns emerge, such as Train Timetabling and Map Labeling appearing close to each other, suggesting potential transfer learning opportunity.
Figure 8: Visualization of MIP embeddings from strIPlib. Forge has never seen these instances, was only trained on 1,800 MIPLIB instances, yet still, manages to cleanly cluster strIPlib problems.
A.4 Ablation Study on The Codebook Size
An important consideration in vector quantization is determining the appropriate codebook size to effectively capture global structural patterns. While increasing the number of codes might intuitively enhance representation capacity, empirical evidence shows this relationship is non-monotonic due to code under-utilization, where some codes remain unused during training. This phenomenon has been extensively studied in domains such as speech and computer vision (Yu et al., 2022; Zeghidour et al., 2021; Lee et al., 2022). To examine this trade-off, we conduct an ablation study measuring NMI scores for clustering across 1,050 D-MIPLIB instances. Following the setup in §4, we generate Forge embeddings with varying codebook sizes.
Table 1 reports our results showing no statistically significant difference in clustering performance, measured by NMI against the ground truth, across different codebook configurations. We attribute this stability to the clustering objective, which appears less sensitive to codebook size variations compared to downstream predictive tasks.This observation aligns with Yang et al. (2024), who reported consistent classification accuracy across various graph datasets until codebook sizes exceeded 16,000, at which point performance began to degrade. Based on these findings, we set the codebook size to 5,000 in our experiments (§5.1 and §5.2). While our pre-trained Forge model was trained on 2,850 MIP instances, making 5,000 codes potentially excessive, the framework is designed as a foundational architecture ready for extension to much larger datasets.
Codebook Size NMI
500 $0.810\pm 0.027$
1,000 $0.818\pm 0.030$
2,500 $0.822\pm 0.026$
5,000 $\textbf{0.843}\pm\textbf{0.031}$
10,000 $0.805\pm 0.022$
Table 1: Ablation study on codebook size of the Forge architecture measuring NMI scores.
A.5 Vector Arithmetic in the Latent MIP Embedding Space
Given that we have embeddings for each MIP instance across various problem types, we ask if we can identify certain ‘directions’ in this latent optimization embedding space that could shift a MIP instance from one problem type to the other based on our theoretical understanding. This line of reasoning is inspired by the earlier work on understanding analogies in word embeddings, as in the famous $King-Man+Woman\approx Queen$ example (Ethayarajh et al., 2018).
Concretely, we inspect the relationship among the following covering and packing problems:
A.5.1 Set Cover Problem (SCP)
Given a universe of elements and a collection of subsets, find the smallest number of subsets that cover all elements.
$min\sum_{S\in\mathcal{S}}x_{S}$
$\sum_{S:e\in S}x_{S}\geq 1\quad\forall e\in\mathcal{U}$
$x_{S}\in\{0,1\}\quad\forall S\in\mathcal{S}$
A.5.2 Vertex Cover Problem
Given a graph, find the smallest set of vertices such that every edge has at least one endpoint in the set.
$min\sum_{v\in V}x_{v}$
$x_{u}+x_{v}\geq 1\quad\forall(u,v)\in E$
$x_{v}\in\{0,1\}\quad\forall v\in V$
A.5.3 Bin Packing Problem
Given a set of items and a collection of bins with a capacity, find the smallest number of bins that pack all items within bin capacities.
$min\sum_{j=1}^{n}y_{j}$
$\sum_{j=1}^{n}x_{ij}=1\quad\forall i\in\{1,\ldots,m\}$
$\sum_{i=1}^{m}w_{i}x_{ij}\leq C\cdot y_{j}\quad\forall j\in\{1,\ldots,n\}$
$x_{ij}\in\{0,1\},\quad y_{j}\in\{0,1\}$
A.5.4 Independent Set Problem
Given a graph, find the largest set of vertices such that no two selected nodes are adjacent.
$max\sum_{v\in V}x_{v}$
$x_{u}+x_{v}\leq 1\quad\forall(u,v)\in E$
$x_{v}\in\{0,1\}\quad\forall v\in V$
Observation: In these problem definitions and formulations, notice how Set Cover and Vertex Cover are similar to each other as covering problems, and Bin Packing and Independent Set are similar to each other as packing problems. Conversely, the pairs are complementary to each other, e.g, the sum of minimum vertex cover and maximum independent set equals to the number of vertices.
Research Question: We ask the following question: if we identify the difference in dimensions between covering and packing, and then push Minimum Vertex Cover instances along that direction, do we obtain embeddings that are closer to Maximal Independent Set instances? The intuition behind this is to remove the ‘cover’ aspect from Vertex Cover, and then add to that the ‘packing’ aspect of Bin Packing, to obtain Independent Set as a result.
To validate this experimentally, we fix the graph size to 1,000 vertices and select 50 random instances each for Set Cover, Vertex Cover, Independent Set from D-MIPLIB (Huang et al., 2024). and 50 random Bin Packing from strIPlib (Bastubbe et al., 2025). We also control for the problem difficulty by ensuring all instances are solvable by Gurobi within 60 seconds.
Figure 9: A (stretch) analogy of $King-Men+Woman\approx Queen$ for Combinatorial Optimization: $Vertex\ Cover-Set\ Cover+Bin\ Packing\approx Independent\ Set$. The updated Vertex Cover instances are shown in black which move closer to the Maximal Independent Set instances.
Vector Arithmetic for Optimization: Next, we apply vector arithmetic on the embeddings of MIP instances to verify our intuition, as a (stretch) analogy to the famous $King-Man+Woman\approx Queen$ example, and test for $VertexCover-SetCover+BinPacking\approx IndependentSet$. We note that this analogy is not perfect, as there is no single ‘King’ instance in the optimization embedding space, but instead, we are averaging over multiple instances while controlling for the same graph size and similar difficulty across problems.
Given the embeddings of instances for Set Cover, $e_{sc}$, Set Packing, $e_{bp}$, Vertex Cover, $e_{mvc}$, and Independent Set, $e_{mis}$, where $e\in\mathbb{R}^{\lvert instances\rvert\times|codes\rvert}$, we compute the difference in dimensions between covering and packing, $d_{sc-bp}$, as follows:
$\mu_{sc}$ $=mean(e_{sc},axis=0)\in\mathbb{R}^{1\times|codes\rvert}$ (5)
$\mu_{bp}$ $=mean(e_{bp},axis=0)\in\mathbb{R}^{1\times|codes\rvert}$ (6)
$d_{sc-bp}$ $=\mu_{sc}-\mu_{bp}\in\mathbb{R}^{1\times|codes\rvert}$ (7)
Given the difference between covering and packing, $d_{sc-bp}$, we update the embedding of the instances of minimum vertex cover problem:
$e_{updated\_mvc}$ $=e_{mvc}-d_{sc-bp}\in\mathbb{R}^{\lvert instances\rvert\times|codes\rvert}$ (8)
Results: Figure 9 visualizes these vector operations in the latent MIP embedding space using PaCMAP in Figure 9-A and PCA in Figure 9-B. In both visualizations, notice how the embeddings of updated Vertex Cover move closer to the embeddings of Independent Set compared to their initial starting point, once they are modified by the direction obtained from the difference of embeddings of covering and packing instances.
Similarly, we further examine the vector arithmetic results for:
  • In Figure 10 above, SetCoverVertexCover+IndependentSet=BinPackingSet\ Cover-Vertex\ Cover+Independent\ Set=Bin\ Packing: Intuitively, this is can be understood as removing the ‘cover’ direction from Set Cover by subtracting Vertex Cover and adding the ‘packing’ direction from Independent Set to push the updated Set Cover instances closer to the Bin Packing instances from the initial Set Cover instances.
  • In Figure 11 below, IndependentSetSetPacking+SetCover=VertexCoverIndependent\ Set-Set\ Packing+Set\ Cover=Vertex\ Cover: Intuitively, this is can be understood as removing the ‘packing’ direction from Independent Set by subtracting from Bin Packing and adding the ‘cover’ direction from Set Cover to push the updated Independent Set instances closer to the Vertex Cover instances from the initial Independent Set instances.
Figure 10: A (stretch) analogy of $King-Man+Woman\approx Queen$ for Combinatorial Optimization: $Set\ Cover-Vertex\ Cover+Independent\ Set\approx Bin\ Packing$. The updated Set Cover instances are shown in black that move closer to the Bin Packing instances.
Figure 11: A (stretch) analogy of $King-Men+Woman\approx Queen$ for Combinatorial Optimization: $Independent\ Set-Bin\ Packing+Set\ Cover\approx Vertex\ Cover$. The updated Independent Set instances are shown in black that move closer to the Minimum Vertex Cover instances.
A.6 Details of Our Experimental Setup
@@
For the experiments presented in §5, we train on the ml.g5.xlarge AWS instance with a GPU with 24 GB of memory. Inference experiments were run on the ml.c5.12xlarge instance with 48 cores and 96 GB of RAM. To ensure consistency and fairness, all experiments were executed with Gurobi (v12.0.3), a state-of-the-art commercial MIP solver (Gurobi Optimization, LLC, 2024), restricted to a single thread and a time limit of 3600 seconds. Unsupervised pre-training and integrality gap fine-tuning are run for 10 epochs with a learning rate of $10^{-4}$, while the search guidance prediction task is trained for 25 epochs using a learning rate of $10^{-5}$. The fixed radius in search guidance is set to 0.1.
A.7 Details of the Triplet Loss
@@
Recall that Forge generates embedding vectors per instance and per variable/constraint. A straightforward method of predicting which variables are likely to be part of the solution is to treat it as a binary classification problem and use binary cross entropy (BCE) loss. However, this poses challenges due to the large class imbalance where most variables are not part of the solution.
@@
Triplet Construction: To address this issue, we generate 5 feasible solutions for each instance. Variables in the instance are grouped by how many of the 5 solutions each variable was repeated in. A variable is considered a ‘negative’ variable if and only if it appears in none of the 5 solutions. Next, we use the standard triplet loss (Schroff et al., 2015) to fine tune the embeddings of the variables. The triplet loss is given by:
@@
Figure 12: PS-Gurobi vs. PS-Gurobi + Forge vs. PS-Gurobi + Random Search Guidance. Each subplot shows primal gap (the lower, the better) averaged across 50 medium instances in each problem.
A.8 Ablation Study on Search Guidance
In §5.2, as part of our comparison with state-of-the-art ML methods, we augmented PS-Gurobi (Han et al., 2023) with Forge embeddings and observed performance gains from their combination. More precisely, in PS-Gurobi, the input node features for constraints and variables are 4 and 6 dimensions, respectively. To minimize architectural changes, we applied Principal Component Analysis (PCA) to reduce the dimensionality of Forge embeddings from 1024 down to 64. This raises an important question: how much of the improvement is due to Forge versus the increased model capacity from higher-dimensional inputs? To control for this effect, we repeated the experiments from §5.2, augmenting PS-Gurobi with 64-dimensional random vectors. This comparison isolates the contribution of our pre-trained embeddings to overall performance. Each PS-Gurobi model for each problem type and augmentation was trained from scratch.
Figure 12 compares PS-Gurobi, PS-Gurobi augmented with Forge, and PS-Gurobi with random vectors. Our ablation shows that adding random vectors occasionally improves performance, likely due to the increase in input dimensionality, from 4 and 6 features for constraints and variables to 68 and 70, resulting in greater model capacity. However, semantic embeddings from Forge consistently dominate, delivering the strongest performance across both problem types.
A.9 Additional Related Work